OpenStudy (superflytyguy):

Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f o g)(-5)

OpenStudy (anonymous):

(f o g) means???

OpenStudy (anonymous):

f o g =f(g(x))

OpenStudy (anonymous):

so what will the answer?

OpenStudy (anonymous):

find g(-5) then the answer of that is x of f(x)

OpenStudy (anonymous):

what's your answer?

OpenStudy (superflytyguy):

I dunno I'm still plugging it in

OpenStudy (anonymous):

...????!!!

OpenStudy (anonymous):

where is the problem could u tell me how do u solve it?

OpenStudy (anonymous):

look f o g=f(g(x)) OK?

OpenStudy (superflytyguy):

Yea

OpenStudy (anonymous):

so just imagine U=g(x) u should find F(U) OK?

OpenStudy (anonymous):

in your question u=g(-5)

OpenStudy (superflytyguy):

thanks :) I'm not that great in math

OpenStudy (anonymous):

yourwelcome U can be great in math if u want